# EvolMPNN: Predicting Mutational Effect on Homologous Proteins by Evolution Encoding

### v1

@@ Appendix B Theoretical Analysis
Theorem 1 (Bourgain Theorem).
@@
Theorem 1 states the Bourgain Theorem [53], which shows the existence of a low distortion embedding that maps from any metric space to the $l_{p}$ metric space.
Theorem 2 (Constructive Proof of Bourgain Theorem).
@@
Then, $f$ is an embedding method that satisfies Theorem 1.
Anchor Selection. EvolMPNN can be viewed as a generalisation of the embedding method of Theorem 2 [54], where $\mathrm{F}_{\textsc{Dist}}(\cdot)$ is generalised via message passing functions (Eq 3-Eq. 6). Therefore, Theorem 2 offers a theoretical guide that $O(\log^{2}M)$ anchors are needed to guarantee low distortion embedding. Following this principle, EvolMPNN choose $k=\log^{2}M$ random anchors, denoted as $\{S_{j}\}_{j=1,2,\dots,\log^{2}M}$, and we sample each protein in $\mathcal{M}$ independently with probability $\frac{1}{2^{j}}$.
Appendix C Datasets
Adeno-associated virus (AAV) capsid proteins are responsible for helping the virus carrying viral DNA into a target cell [61]; there is great interest in engineering versions of these proteins for gene therapy [14, 62, 63]. [14] produces mutants on a 28 amino-acid window from position 561 to 588 of VP-1 and measures the fitness of resulting variants with between 1 and 39 mutations. We adopt three splits from the benchmark [4], including 2-vs-Rest, 7-vs-Rest and Low-vs-High.
GB1 is the binding domain of protein G, an immunoglobulin binding protein found in Streptococcal bacteria [64, 65]. [16] measure the fitness of generated mutations. We adopt three splits from the benchmark [4], including 2-vs-Rest, 3-vs-Rest and Low-vs-High.
The green fluorescent protein is an important marker protein, enabling scientists to see the presence of the particular protein in an organic structure by its green fluorescence [66]. [15] assess the fitness of green fluorescent protein mutants. We adopt one available split, 3-vs-Rest, from the benchmark [5]. Besides, in order to evaluate the models’ effectiveness, we add two splits, 2-vs-Rest and Low-vs-High, in this paper.
Appendix D Baselines
Table 4: Description and implementation of baseline methods.
Method Descriotion Encoder Pooling Output layer
Levenshtein
Levenshtein distance to wild
-type.
BLOSUM62
BLOSUM62-score relative
to wild-type.
DDE
ipeptide Deviation from
Expected Mean
2-layer MLP - -
Moran Moran correlation 2-layer MLP - -
LSTM Simple LSTM model 3-layer LSTM Weighted sum pool 2-layer MLP
Transformer Simple Transformer model
4-layer Transformer,
4 attention heads
- 2-layer MLP
CNN Simple convolutional model 2-layer CNN Max pool 2-layer MLP
ResNet
Classic framework of skip co-
nnections and residual blocks
8-layer ResNet
Attentive weighted sum 2-layer MLP
ProtBert
750M param transformer
pre-trained on Uniref 50
30-layer BERT
16 attention heads
Linear pool 2-layer MLP
ESM-1b
650M param transformer
pre-trained on Uniref 50
- Mean pool 2-layer MLP
ESM-1v
650M param transformer
pre-trained on Uniref 90
- Mean pool 2-layer MLP
ESM-2
3B param transformer
pre-trained on Uniref 50
- Mean pool 2-layer MLP
GCN Graph convolutional network
2/3-layer GCN encoder
- 2-layer MLP
GAT Graph attention network
2/3-layer GAT encoder
- 2-layer MLP
GraphTransf.
Transformer model designed
for graphs
2/3-layer GraphTransf.
encoder
- 1-layer MLP
Nodeformer
Efficient Transformer variant
design for graphs
2/3/4-layer Nodeformer
encoder
- 1-layer MLP
EvolMPNN -
2/3-layer Residue encoder
and Evolution encoder
- 1-layer MLP
EvolGNN -
2/3-layer Residue encoder
and Evolution encoder
- 1-layer MLP
EvolFormer -
2/3-layer Residue encoder
and Evolution encoder
- 1-layer MLP
We present details about our baseline and proposed models in Table 4.

### v2

@@ Appendix 0.A Theoretical Analysis
Theorem 0.A.1 (Bourgain Theorem)
@@
Theorem 0.A.1 states the Bourgain Theorem [3], which shows the existence of a low distortion embedding that maps from any metric space to the $l_{p}$ metric space.
Theorem 0.A.2 (Constructive Proof of Bourgain Theorem)
@@
Then, $f$ is an embedding method that satisfies Theorem 0.A.1.
Anchor Selection. EvolMPNN can be viewed as a generalisation of the embedding method of Theorem 0.A.2 [24], where $\mathrm{F}_{\textsc{Dist}}(\cdot)$ is generalised via message passing functions (Eq 3-Eq. 6). Therefore, Theorem 0.A.2 offers a theoretical guide that $O(\log^{2}M)$ anchors are needed to guarantee low distortion embedding. Following this principle, EvolMPNN choose $k=\log^{2}M$ random anchors, denoted as $\{S_{j}\}_{j=1,2,\dots,\log^{2}M}$, and we sample each protein in $\mathcal{M}$ independently with probability $\frac{1}{2^{j}}$.